OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) equals 4/x. and g(x) equals 4/x

hero (hero):

Hint: \[f(x) = \frac{4}{x}\] \[f(g(x)) = \frac{4}{g(x)}\]

OpenStudy (anonymous):

uh... Im sorry. I still dont really understand

OpenStudy (anonymous):

g(x) = 4/x, so wouldnt what you put be like, 4/4/x/4/x????

hero (hero):

Actually, \(\dfrac{4}{\dfrac{4}{x}} = 4 \div \dfrac{4}{x} = 4 \times \dfrac{x}{4} = x\)

hero (hero):

@mihirb

OpenStudy (anonymous):

oh srry @Hero